OpenStudy (anonymous):

The measures of three sides of a triangle are consecutive odd integers. The perimeter of the triangle is 57. What are the measures of the three sides? One point for the equation and one point for the solution

OpenStudy (nali):

consecutive odd integers can be represented like the following n n+2 n+4 so forexample if n= 1 it would be 1,3,5 if added with 2 or 4 right

OpenStudy (nali):

so the perimeter of any triangle is the addtion of all the 3 sides therefore n+(n+2)+(n+4)=57 3n+6=57 3n=51 and then just solve for n and whatever it equals just plug in the other value n, n+2, n+4 and u will get ur three sides